OpenStudy (anonymous):

hi all can you give me of method s=1+2+3+.........+n ?

OpenStudy (joannablackwelder):

Summation notation

OpenStudy (anonymous):

If you would like to sum such a series of numbers, you first determine if the elements are an even or an odd number. In case we have an even number of elements, we add the first and the last element and multiply by half the number of the elements. Example: \[S = 1+2+3+4+5+6 = (6+1)*6/2 =7*3 = 21\] If the number of elements in the series is odd, you ignore the one in the middle and repeat the same procedure. After you've calculated their sum, you simply add the middle element. Example: \[S=3+4+5+6+7 = 10*4/2+5=25\] This is probablly the easiest practical method for summation developped by Gauss. If, after all, you need a general formula, where the elements can be any real number, you need the sum of an arithmetic progression: \[s=\frac{ n*(a _{1}+a _{n}) }{ 2 }\] where n is the nu mber of elements, a1 is the smallest one and an - the largest one. Example: \[S = 3 + 7 +9+10 = \frac{ 4*(3+10) }{ 2 }=26\]

OpenStudy (anonymous):

can i ask you some question ?

OpenStudy (anonymous):

@khemvatey what is it? I can try answering

OpenStudy (anonymous):

S(n)=1+3+............+2n-1 and 1*2*3*...........*n